# Computing the Elementary Symmetric Polynomials in Positive Characteristics

### v2

@@ 3.3 Border Depth-Three Formulas
Now, in this section, we will connect the border symmetric computational model to $\overline{\Sigma^{[k]}\Pi\Sigma}$ formulas. In [Kum20], the main result came from observing we could convert a slightly restrictive extension of the border symmetric model to $\overline{\Sigma^{[2]}\Pi\Sigma}$ formulas. Specifically, for a given homogeneous degree $d$ $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ such that there are linear $L_{1},\ldots,L_{m}\in\mathbb{F}(\epsilon)[x_{1},\ldots,x_{n}]$ where $e_{d}^{m}(L_{1},\ldots,L_{m})\simeq f$ and $e_{k}^{m}(L_{1},\ldots,L_{m})=0$ for every $k<d$, then we can write $f$ as a $\overline{\Sigma^{[2]}\Pi\Sigma}$ formula, using (1),
@@
To prove this, let $f_{1}^{(1)},\ldots,f_{m_{k}}^{(k)}\in\mathbb{F}(\epsilon)[x_{1},\ldots,x_{n}]$ be affine and $c_{1},\ldots,c_{k}\in\mathbb{F}(\epsilon)$ be such that $f\simeq\sum_{i=1}^{k}c_{i}\prod_{j=1}^{m_{i}}f_{j}^{(i)}$. We claim that, if each $f_{j}^{(i)}$ is not constant-free, then the result is obvious. Notice that, in this case, we can assume that each $f_{j}^{(i)}(0)=1$. Then, observe that
$f$ $=H_{d}[f]\simeq H_{d}\left[\sum_{i=1}^{k}c_{i}\prod_{j=1}^{m_{i}}L_{j}^{(i)}\right]=\sum_{i=1}^{k}c_{i}\cdot H_{d}\left[\prod_{j=1}^{m_{i}}L_{j}^{(i)}\right]=\sum_{i=1}^{k}c_{i}\cdot e_{d}(L_{1}^{(i)},\ldots,L_{m_{i}}^{(i)}).$
Now, we only need to consider what happens if there is some $f_{j}^{(i)}$ such that $f_{j}^{(i)}(0)=0$. We will show that we can modify this polynomial to make it have a constant part without changing the polynomial we approximate. To do this, we merely need to add a constant part to $f_{j}^{(i)}$ that will not change that polynomial we approximate. The proof is then completed from the following lemma.
@@
To prove this, imagine that $\alpha$ is a new independent variable. Then, consider $F\cdot(\ell+\alpha)+G$ as a polynomial over $x_{1},\ldots,x_{n},\alpha$. Notice that the coefficients of monomials in $x_{1},\ldots,x_{n}$ are in $\mathbb{F}[\epsilon]$. Then, observe that there are only a finite number of coefficients in monomials that include $\alpha$. Therefore, it is obvious that we can select $\alpha$ to make all of these coefficients polynomials (multiplying the denominators) and, by multiplying by $\epsilon^{N}$ (for some large enough $N$), we ensure that this does not change the approximation. ∎
@@ 4 Future Directions
  • In Lemma 1.7, we observe that there are cases where $\dim V_{2}(e_{d}^{n})=d-2$ and other cases where $\dim V_{2}(e_{d}^{n})=d-1$, but these cases do not cover all possible ones. Can we determine a characterization for, given a $\operatorname{char}(\mathbb{F})$, $d$, and $n$, which case $e_{d}^{n}$ falls under?
  • When considering Theorem 1.2, we, motivated by [Shp02], force the inputs to the elementary symmetric polynomials to be linear. We could then, just as easily, allow the inputs to be affine and then increase the degree of the elementary symmetric polynomial. Of course, we can easily homogenize this case by introducing a new independent variable $x_{0}$, so this case can be considered as an extension of the case with linear inputs. Unfortunately, due to limitations with Newton’s identities, our current strategy only lets us consider degrees of $\operatorname{char}(\mathbb{F})+1$. Is there a way to show that a particular polynomial cannot be computed by this inhomogeneous model?
@@ Acknowledgements
Thank you to my advisor Dr. Srikanth Srinivasan for encouraging and helping me throughout this entire line of research and steering me in the right directions through regular chats. A lot of the ideas in this paper came from our long discussions. Also, thank you for helping during the process of writing this paper, where your extensive knowledge of the field was paramount. Further, thank you to Theo Fabris for being involved in some of these aforementioned discussions.
@@ Appendix B Formula Lower Bounds
In this section, we will focus on proving the results of [CKSV22], specifically, we will prove Lemma 2.1. This is done for the sake of completeness and because the statement that we use, while not difficult to see from the proof in the original paper, is slightly different. Further, we provide a minor simplification to the original proof.
@@
First notice that the formal degree of a formula monotonically increases as one goes from a leaf to the root. The leaves of the formula have formal degree at most one and the root has formal degree $d$. Then, we observe that the formula degree of an internal node is at most the sum of the formal degrees of its children. It is, therefore, impossible for a parent node to have formal degree above 2t12t-1 with children of formal degree strictly less than $t$.

### v3

@@ 3.3 Border Depth-Three Formulas
Now, in this section, we will connect the border symmetric computational model to $\overline{\Sigma^{[k]}\Pi\Sigma}$ formulas. In [22], the main result came from observing we could convert a slightly restrictive extension of the border symmetric model to $\overline{\Sigma^{[2]}\Pi\Sigma}$ formulas. Specifically, for a given homogeneous degree $d$ $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ such that there are linear $L_{1},\ldots,L_{m}\in\mathbb{F}(\epsilon)[x_{1},\ldots,x_{n}]$ where $e_{d}^{m}(L_{1},\ldots,L_{m})\simeq f$ [^6] Meaning that it approximates in the border setting and $e_{k}^{m}(L_{1},\ldots,L_{m})=0$ for every $k<d$, then we can write $f$ as a $\overline{\Sigma^{[2]}\Pi\Sigma}$ formula, using (1),
@@
To prove this, let $f_{1}^{(1)},\ldots,f_{m_{k}}^{(k)}\in\mathbb{F}(\epsilon)[x_{1},\ldots,x_{n}]$ be affine and $c_{1},\ldots,c_{k}\in\mathbb{F}(\epsilon)$ be such that $f\simeq\sum_{i=1}^{k}c_{i}\prod_{j=1}^{m_{i}}f_{j}^{(i)}$. We claim that, if each $f_{j}^{(i)}$ is not constant-free, then the result is obvious. Notice that, in this case, we can assume that each $f_{j}^{(i)}(0)=1$ by dividing out each affine functions constant term. Then, observe that
$f$ $=H_{d}[f]\simeq H_{d}\left[\sum_{i=1}^{k}c_{i}\prod_{j=1}^{m_{i}}f_{j}^{(i)}\right]=\sum_{i=1}^{k}c_{i}\cdot H_{d}\left[\prod_{j=1}^{m_{i}}f_{j}^{(i)}\right]=\sum_{i=1}^{k}c_{i}\cdot e_{d}^{m_{i}}(f_{1}^{(i)},\ldots,f_{m_{i}}^{(i)}).$
Now, we only need to consider what happens if there is some $f_{j}^{(i)}$ such that $f_{j}^{(i)}(0)=0$. We will show that we can modify this polynomial to make it have a constant part without changing the polynomial we approximate. The proof is then completed from the following lemma.
@@
To prove this, imagine that $\alpha$ is a new independent variable. Then, consider $F\cdot(\ell+\alpha)+G$ as a polynomial over $x_{1},\ldots,x_{n},\alpha$. Notice that the coefficients of monomials in $x_{1},\ldots,x_{n}$ are in $\mathbb{F}(\epsilon)[\alpha]$. Then, observe that there are only a finite number of coefficients in monomials that include $\alpha$. Therefore, it is obvious that we can select $\alpha$ to make all of these coefficients polynomials (multiplying the denominators) and, by multiplying this $\alpha$ value by $\epsilon^{N}$ (for some large enough $N$), we ensure that this does not change the approximation. ∎
@@ 4 Future Directions
  • When considering Theorem 1.2, we, motivated by [29], analyzed projections of the elementary symmetric polynomials by linear functions [^7] This is done in their case as linear functions are sufficient to make the model universal. It would be natural to relax this model to allow projections of affine functions. Of course, we can easily homogenize this case by introducing a new independent variable $x_{0}$, so this case can be considered as an extension of the case with linear inputs. We then, unfortunately, face the issue that this homogenization could increase the degree of the polynomial, which is a problem for our results, as they only work with degrees equal to $\operatorname{char}(\mathbb{F})+1$. Can we extend our characterizations to work for polynomials of arbitrary degrees?
  • Recall that the border affine Chow rank of a polynomial $f$ is given by the smallest $k$ such that $f$ can be computed by $\overline{\Sigma^{[k]}\Pi\Sigma}$. When the characteristic is zero, it is known that this value is at most two, but we have now shown that these results do not extend to positive characteristics (we get lower bounds linear in the number of variables). It would be an interesting line of research to study what sorts of upper bounds we can find when the characteristic is positive.
@@ Acknowledgements
Thank you to my advisor Srikanth Srinivasan for encouraging and helping me throughout this entire line of research and steering me in the right directions through regular chats. A lot of the ideas in this paper came from our long discussions. Also, thank you for helping during the process of writing this paper, where your extensive knowledge of the field was paramount. Further, thank you to Theo Fabris for being involved in some of these aforementioned discussions.
@@ Appendix B Formula Lower Bounds
In this section, we will focus on proving the results of [6], specifically, we will prove Lemma 2.1. This is done for the sake of completeness and because the statement that we use, while not difficult to see from the proof in the original paper, is slightly different. Further, we provide a minor simplification to the original proof.
@@
First notice that the formal degree of a formula monotonically increases as one goes from a leaf to root. The leaves of the formula have formal degree at most one and the root has formal degree $d$. Then, we observe that the formal degree of an internal node is at most the sum of the formal degrees of its children. It is, therefore, impossible for a parent node to have formal degree above 2t12t-1 with children of formal degree strictly less than $t$.